QUESTION 2: COMPANY EQUITY, SPECIAL REPORTING ISSUES AND EPS Question 2 consists of three separate parts, Part 1, Part 2 and Please show all working
Answer #1

In the books of Kelly Ltd.

Journal (in $)

DATE (2018) PARTICULARS DEBIT CREDIT
April 1

Share Issue expenses A/C Dr.

To Underwriters (Equity) A/C

To Underwriters (Preference) A/C

(On being underwriting commission made due which is to be paid to underwriters.)

4500

500

4000

April 10

Bank A/C Dr.

To Share Application (Equity) A/C

To Share Application (Preference) A/C

(On being money received on share application.)

380000

210000

170000

April 15

Share Application (Equity) A/C Dr.

Share Application (Preference) A/C Dr.

Underwriters (Preference) A/C Dr.

To Equity Share capital A/C

To Preference Share capital A/C

To Bank A/C

(On being ordinary and preference shares being allotted and money on rejected ordinary shares application being refunded through bank.)

210000

170000

30000

150000

200000

60000

April 20

Bank A/C Dr.

To Underwriters (Preference) A/C

(On being amount for preference shares issued to underwriters received less underwriting commission.)

26000

26000

orking notes:

  • Amount to be received towards application for shares. Ordinary shares (including 40,000 oversubscribed) =>140000 * 1.5 = $2,10,000. Preference shares (excluding 15,000 not subscribed) => 85000 * 2 = $1,70,000
  • Underwriters to be debited for issue of preference share capital will be for 15,000 preference shares of $2 each so by $30,000.
  • Amount to be received by underwriters of preference shares towards shares issued to them will be money to be received towards issue of preference shares less underwriting commission for preference shares ($4,000), i.e.,30000-4000 = $26,000
  • Amount refunded through bank for rejected applications of Ordinary shares will be of 40,000 applications extra received with $1.5 each. Amount to be refunded = 40000 * 1.5 = $60,000